The average of 7 numbers is 39. Two of the numbers whose sum is 83 are removed. What is the average of the remaining numbers?

Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)


To calculate the average of 7 numbers, you must have added the seven numbers and then divided by 7. So if that result is 39, the sum of the seven numbers must have been 7 times 39. From that amount, subtract 83. Then, since you removed two numbers, divide by 5.

John


Answer by Alan3354(69443)

The average of 7 numbers is 39. Two of the numbers whose sum is 83 are removed. What is the average of the remaining numbers?
--------------
The sum is the avg x the # of members = 39*7 = 273
(273 - 83) = 190
There are now 5 numbers, so the avg is 190/5 = 38